How much pure acid should be mixed with 7 gallons of a 40% acid solution in order to get an 80% acid solution?
Answer by stanbon(29561)

How much pure acid should be mixed with 7 gallons of a 40% acid solution in order to get an 80% acid solution?
------------
acid + acid = acid
0.40*7 + 1.00x = 0.80(7+x)
40*7 + 100x = 80(7+x)
280 + 100x = 560 80x
20x = 280
x = 14 gallons (amount of 100% acid that must be added)
============
Cheers,
Stan H.